# UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

### v3

@@ Abstract
Evaluating the open-ended outputs of Large Multimodal Models has become a bottleneck as model capabilities, task diversity, and modality rapidly expand. Existing "LLM-as-a-Judge" evaluators are typically narrow in specific tasks and aspects. In this paper, we argue that, on one hand, based on the interconnected nature of aspects, learning specific aspects can generalize to unseen aspects; on the other hand, jointly learning to assess multiple visual aspects and tasks may foster a synergistic effect. To this end, we propose UFEval, the first unified fine-grained evaluator with task and aspect generalization for four evaluation tasks — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. Specifically, (1) We first construct a hierarchical aspect taxonomy encompassing 112 distinct aspects across the aforementioned four tasks. (2) Then, building upon this taxonomy, we create FRABench, a fine-grained evaluation dataset comprising 60.4k pairwise samples with 325k evaluation labels obtained from a combination of human and GPT-4o annotations. FRABench provides a large-scale, multi-modal, and aspect-level resource for training and testing evaluators. (3) Finally, leveraging FRABench, we develop UFEval, a unified fine-grained evaluator. Experiments show that learning on specific aspects enables UFEval to generalize to unseen aspects, and joint learning to assess diverse tasks and aspects can lead to substantial mutual benefits.
Codehttps://github.com/ALEX-nlp/UFEval
Datasetshttps://huggingface.co/datasets/SPUH/FRABench
Introduction
@@
As Large Multimodal Models have shown amazing abilities in human-like question answering (Grattafiori et al. 2024; Wang et al. 2024a), assessing the quality of their free-form outputs has become increasingly challenging. Automated evaluators, a.k.a. “LLM-as-a-Judge” paradigm (Saha et al. 2025) have therefore received much research attention (Li et al. 2024b). Despite their progress, we posit two concerns: (1) Current evaluators (Liu et al. 2023b; Ke et al. 2023) are typically tailored to specific aspects, which limits their adaptability to unseen aspects. (2) They are also limited to specific tasks and modalities, which sharply constrains their scope of application. Table 1 shows the detailed comparison with existing evaluators. We intuitively argue that, on one hand, evaluation aspects are inherently interconnected (Fu et al. 2023). Specifically, aspects such as engagement, naturalness, and creativity are closely linked. Thus, similar semantics and evaluation standards can be transferred across diverse tasks. On the other hand, jointly learning to assess multiple visual aspects and tasks may foster synergistic effects (Wang et al. 2025). For instance, learning object alignment in image captioning improves character consistency evaluation in multi-image scenarios through better object recognition. Furthermore, progress in image understanding tasks can enhance the evaluation of image generation results by enabling more accurate judgments of content quality and contextual appropriateness. This cross-aspect and cross-task synergy enables a more robust evaluation for diverse outputs. It inspires the development of a unified, fine-grained evaluator to achieve better aspect of generalization and evaluation performance.
Method Task Modalty
@@
Auto-J
X-Eval
Prometheus 2
Themis
LLaVA-Critic
ImageReward
VisionReward
UFEval (Ours)
Table 1: Comparison of our UFEval with recent representative evaluators, including Auto-J (Li et al. 2023b), X-Eval (Liu et al. 2023b), Prometheus 2 (Kim et al. 2024a), Themis (Hu et al. 2024b), LLaVA-Critic (Xiong et al. 2024), ImageReward (Xu et al. 2023a), and VisionReward (Xu et al. 2024). NLG, IU, IG, and ITIG represent Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation, respectively.
@@
To this end, we propose UFEval, the first unified fine-grained evaluator with task and aspect generalization for four evaluation tasks (i.e., NLG, IU, IG, and ITIG) across 28 sub-tasks. However, training such a unified evaluator requires large-scale, multi-modal, aspect-level evaluation datasets, which are currently unavailable. Therefore, we develop the Fine-grained Aspect Benchmark (FRABench) to address this gap. Specifically, we first conduct a survey on the four tasks to identify key evaluation aspects. We then manually organize, extend, and redefine 112 distinct aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Leveraging the aspect tree, we construct FRABench to ensure comprehensive aspect coverage. We leverage the aspect tree to select relevant multiple aspects for each of the 60.4k pairwise responses. To obtain high-quality evaluation labels while maintaining cost efficiency, we collect available human annotations wherever possible and supplement them with GPT-4o-assisted completions. In total, this process yields 325k evaluation labels. Using FRABench, we develop UFEval, a unified and fine-grained evaluator.
@@
Our experiments demonstrate that UFEval exhibits excellent evaluation quality and strong task and aspect generalization capabilities. This is attributed to learning multiple aspects and tasks jointly, which yields significant mutual enhancement, as well as to fully leveraging the interconnections between learning and unlearning aspects. By conducting ablation experiments across diverse aspects and task baselines, we observe a gradual improvement in evaluation quality, thereby validating our hypothesis.
@@
In summary, our contributions are as follows: (1) We construct FRABench, the large-scale multi-modal aspect-level evaluation dataset. (2) We develop UFEval, the first unified fine-grained evaluator for multiple tasks assessment with task and aspect generalization. (3) Our experiments show that inter-aspect correlations enable generalizable capability, and learning to assess multiple tasks and aspects jointly leads to a synergistic improvement in evaluation performance.
Related Work
Coarse-grained Specific-task Evaluation
@@
Previously, coarse-grained evaluators, which produce overall judgments based on one or several aspects, have been widely explored across various evaluation tasks (Zhu, Wang, and Wang 2023; Xu et al. 2023b; Ye et al. 2023; Liu et al. 2024b; Wang et al. 2023a; Jiang et al. 2023). For instance, in NLG, Wang et al. (Wang et al. 2023b) first introduce PandaLM, a fine-tuned LLM designed to evaluate pairwise texts based on several aspect. Similarly, Li et al. (Li et al. 2023b) propose Auto-J, an evaluator capable of handling a broader range of tasks and aspects, supporting both pointwise and pairwise evaluation settings. In image generation evaluation, Xu et al. (Xu et al. 2023a) develop ImageReward, an evaluator trained on a large-scale human annotation dataset. Using alignment and fidelity as reference dimensions, ImageReward provides an overall assessment. Coarse-grained evaluations, while effective overall, lack the granularity to identify specific model weaknesses and often introduce aspect bias.
Fine-grained Specific-task Evaluation
@@
To address the limitations of coarse-grained evaluation, recent studies have shifted toward fine-grained evaluation (Ye et al. 2023; Kim et al. 2024b; Ke et al. 2023; Kim et al. 2023; Li et al. 2023d; Ying et al. 2024; Bai et al. 2023), where LMs are fine-tuned on multi-aspect datasets to produce aspect-specific judgments. For example, in NLG, Hu et al.(Hu et al. 2024b) propose Themis, an LLM trained on the GPT-4 annotated NLG-Eval corpus. In IU, LLaVA-Critic(Xiong et al. 2024) is the first fine-grained evaluator integrating diverse criteria, showing strong correlation with GPT-4o. For IG, Xu et al.(Xu et al. 2024) introduce VisionReward, a VQA-based evaluator that assesses image quality across fine-grained aspects. While these methods improve over coarse-grained evaluation by targeting specific aspects, they struggle with scalability across different tasks and aspects. Liu et al.(Liu et al. 2023b) further investigate aspect generalization through X-Eval, a two-stage learning framework that incorporates auxiliary aspects and demonstrates generalization in NLG. However, its reliance on predefined reference aspects and lack of open-source implementation restricts its usability and reproducibility.
Methodology
@@
To develop UFEval, we need to construct a large-scale, multi-modal, and aspect-level evaluation dataset, which is currently unavailable. The construction pipeline involves two main steps: (1) Evaluation Aspect Construction, and (2) Fine-grained Evaluation Dataset Construction. The following sections detail these two steps.
Evaluation Aspect Construction
Aspect Collection and Extension.
@@
Figure 1: An illustration of the evaluation pipeline. The pipeline consists of two steps: (1) Aspect Selection: First, appropriate aspects are selected from the TAs and UAs Trees based on task property and output modality. As illustrated, the question focuses on “cat image next to the orange dog picture”, while the image shows no cat adjacent to the orange dog. This discrepancy enables evaluation of the model’s ability to avoid hallucinations, specifically its performance on Context Inconsistency. Second, given the text output modality, aspects are selected from the Text Branch in the UAs Tree. (2) Evaluating: UFEval generates comprehensive feedback and relative scores based on the input content and selected aspects.
Aspect Taxonomy Construction.
@@
  • Universal Aspects (UAs): These aspects exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image).
  • Task-specific Aspects (TAs): These aspects exhibit task dependence, as they primarily focus on task property rather than output quality. For example, engagingness for story generation or accuracy for mathematical reasoning.
Source Tasks Dataset Sub-Tasks Split Size
Public NLG $\blacksquare$ Auto-J Summarization, Creative Writing, Rewriting, etc. Train / FRA-ID 2.5k / 0.6k
@@
Generated IU $\blacksquare$ TextVQA Text Reasoning Train / FRA-ID 2.1k / 0.4 k
@@
Figure 2: Data statistics of the FRAbench. FRABench comprises the largest number of evaluation aspects among existing aspect-level evaluation datasets. Detailed counts of evaluation labels for each sub-task are in Appendix B.
For hierarchical construction within the two subtrees, we directly adopt existing aspect tree structures from previous studies (Hu et al. 2024a; Ghosh, Hajishirzi, and Schmidt 2023; Huang et al. 2023; DeAlcala et al. 2023; Zeng et al. 2023; Huang et al. 2025). For aspects lacking predefined hierarchies, we determine their positions within the tree based on their definitions—either assigning them to locations within existing structures or establishing them as leaf nodes of root node. Based on the aspect tree, the fine-grained evaluation process can be streamlined, as illustrated in Figure 1.
Fine-grained Evaluation Dataset Construction
Based on the aspect tree, we can construct fine-grained evaluation datasets for training and testing UFEval. Specifically, we construct FRABench—a large-scale, multi-modal, and aspect-level evaluation dataset. Following prior studies (Kim et al. 2024a; Ye et al. 2024) that demonstrate pointwise scoring is more susceptible to contextual bias and available for reward model training, we adopt a pairwise comparison evaluation method in FRABench.
FRABench Construction.
We first collect questions from the datasets corresponding to the 28 sub-tasks we gathered, and generate pairwise samples. Specifically, we first obtain 29.3K response pairs from public datasets and generate the remaining 30.1K pairs using LMMs (model details provided in Appendix B). After obtaining the queries and response pairs, we assign aspects from our aspects tree. As a result, each pairwise sample is annotated with an average of 8 UAs and 3 TAs (Aspect selection information in Appendix B). We then generate evaluation labels via two approaches: (1) Human annotations: directly using three-aspect human-annotated scores from ImageRewardDB (Xu et al. 2023a), supplemented with feedback generated by GPT-4o; (2) GPT-4o annotations: due to the lack of human annotations for most aspects, we use GPT-4o to generate evaluation labels for all other pairwise samples. All templates are provided in Appendix F. Moreover, to mitigate position bias (Ye et al. 2024), we balance the number of samples where response 1 is preferred over response 2 and vice versa by reversing the response positions in half of the surplus samples from the majority class and re-annotating them accordingly. Finally, we generate 325k fine-grained evaluation labels.
Dataset Partitioning and Evaluator Training.
@@
Due to the lack of a large-scale aspect-level meta-benchmark to verify UFEval’s aspect generalizable capability and support our arguments, we can only divide FRABench into training, In-Domain test set (FRA-ID), and Out-of-Domain test set (FRA-OOD). The training and FRA-ID consist of 18 randomly selected sub-tasks from four tasks, covering 22 UAs and 35 TAs, while FRA-OOD contains 10 unseen sub-tasks with 28 seen UAs and 27 unseen TAs. Finally, the training set contains 255.4K samples, FRA-ID contains 45.2K samples, and FRA-OOD contains 24.4K samples. Detailed statistics are presented in Figure 2 and Appendix B.
@@
The role of an evaluator is to assess model responses in place of humans. Therefore, the primary criterion for evaluating an evaluator’s effectiveness is their alignment with human judgments. To this end, we construct corresponding evaluation datasets to measure the consistency. We extracted partial samples from FRA-ID and FRA-OOD, and recruited three humans with master’s degrees for annotation. Finally, we retain 6.9K in-domain evaluation samples (FRA-ID-H) and 6.0K out-of-domain evaluation samples (FRA-OOD-H), respectively. Details of the human annotation process and annotation consistency are provided in Appendix E.
When training UFEval, we use SFT to fine-tune Qwen2-VL-7B-Instruct on the training set. The detailed training configurations and methods are presented in Appendix C.
Experiments

### v4

# FRABench and UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

@@ Abstract
Evaluating open-ended outputs of Multimodal Large Language Models has become a bottleneck as model capabilities, task diversity, and modality rapidly expand. Existing “MLLM-as-a-Judge” evaluators, though promising, remain constrained to specific tasks and aspects. In this paper, we argue that, on one hand, based on the interconnected nature of aspects, learning specific aspects can generalize to unseen aspects; on the other hand, jointly learning to assess multiple visual aspects and tasks may foster a synergistic effect. To this end, we propose UFEval, the first unified fine-grained evaluator with task and aspect generalization for four evaluation tasks — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. However, training such a unified evaluator is hindered by the lack of a large-scale, multi-modal, and aspect-level resource. To address this gap, we introduce FRABench, a comprehensive fine-grained evaluation dataset. Specifically, (1) We first construct a hierarchical aspect taxonomy encompassing 112 distinct aspects across the aforementioned four tasks. (2) Based on this taxonomy, we create FRABench, comprising 60.4k pairwise samples with 325k evaluation labels obtained from a combination of human and GPT-4o annotations. (3) Finally, leveraging FRABench, we develop UFEval, a unified fine-grained evaluator. Experiments show that learning on specific aspects enables UFEval to generalize to unseen aspects, and joint learning to assess diverse visual tasks and aspects can lead to substantial mutual benefits [^1] Our code and dataset can be found at https://github.com/ALEX-nlp/UFEval.
1 Introduction
@@
As Multimodal Large Language Models (MLLMs) have shown amazing abilities in human‑like question answering (Grattafiori et al., 2024), assessing the quality of their free‑form outputs has become increasingly challenging. Automated evaluators, a.k.a. “MLLM‑as‑a‑Judge” paradigm (Saha et al., 2025) have therefore received much research attention (Li et al., 2024c). Despite their progress, we posit two concerns: (1) Current evaluators (Liu et al., 2023b; Ke et al., 2023) are typically tailored to specific aspects, which limits their adaptability to unseen aspects. (2) They are also limited to specific tasks and modalities, which sharply constrains their scope of application. Table 1 shows the detailed comparison with existing evaluators. We intuitively argue that, on one hand, evaluation aspects are inherently interconnected (Fu et al., 2023). Specifically, aspects such as engagement, naturalness, and creativity are closely linked. Thus, similar semantics and evaluation standards can be transferred across diverse tasks. On the other hand, jointly learning to assess multiple visual aspects and tasks may foster synergistic effects (Wang et al., 2025). For instance, learning object alignment in image captioning improves character consistency evaluation in multi-image scenarios, while progress in image understanding enhances image generation evaluation through better judgments of content quality and contextual appropriateness. This cross-aspect and cross-task synergy motivates the development of a unified, fine-grained evaluator for improved generalization and performance.
Table 1: Comparison of our UFEval with recent evaluators. NLG, IU, IG, and ITIG represent Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation, respectively. “–” indicates that the number of supported aspects is not explicitly specified.
Method Task Modalty Aspect Generalizable
@@
AUTO-J (Li et al., 2023b) 332
X-Eval (Liu et al., 2023b) 27
Prometheus 2 (Kim et al., 2024a) -
Themis (Hu et al., 2024b) 50
ImageReward (Xu et al., 2023a) 3
VisionReward (Xu et al., 2024) 37
LLaVA-Critic (Xiong et al., 2024)
UFEval (ours) 112
@@
To this end, we propose UFEval, the first unified fine-grained evaluator with task and aspect generalization for four evaluation tasks (i.e., NLG, IU, IG, and ITIG) across 28 sub-tasks. However, training such a unified evaluator requires large-scale, multi-modal, aspect-level evaluation datasets, which are currently unavailable. Therefore, we develop the Fine-grained Aspect Benchmark (FRABench) to address this gap. Specifically, we first conduct a survey on the four tasks to identify key evaluation aspects. We then manually organize, extend, and redefine 112 distinct aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Using this aspect tree, we construct FRABench with comprehensive aspect coverage. The aspect tree guides us in selecting multiple relevant aspects for each of the 60.4k pairwise responses, for which we obtain evaluation labels through a hybrid approach combining available human annotations with GPT-4o-assisted completions, ensuring cost efficiency. In total, this process yields 325k evaluation labels. Using FRABench, we develop UFEval, a unified and fine-grained evaluator.
@@
Our experiments show that UFEval exhibits excellent evaluation quality and aspect generalization capabilities. This is attributed to learning multiple aspects and tasks jointly, which yields significant mutual enhancement, as well as to fully leveraging the interconnections between learning and unlearning aspects. By conducting ablation experiments across diverse aspects and task baselines, we observe a gradual improvement in evaluation quality, thereby validating our hypothesis. Additionally, we demonstrate the value of using UFEval for preference data generation to align the outputs of models with human preferences via direct preference optimization (DPO). These results validate the effectiveness of the FRABench as a valuable resource for training unified evaluators.
@@
In summary, our contributions are as follows: (1) We construct FRABench, a large-scale multi-modal aspect-level evaluation dataset to train and test evaluators. (2) Upon FRABench, we develop UFEval, the first unified fine-grained evaluator for multiple tasks assessment with task and aspect generalization. (3) Our experiments show that inter-aspect correlations enable generalizable capability, and learning to assess multiple visual tasks and aspects jointly leads to a synergistic improvement in evaluation performance, while also demonstrating UFEval’s value for preference alignment.
2 Related Work
2.1 Coarse-grained Specific-task Evaluation
@@
Previously, coarse-grained evaluators, which produce overall judgments based on one or several aspects, have been widely explored across various evaluation tasks (Zhu et al., 2023; Xu et al., 2023b; Ye et al., 2023; Liu et al., 2024b; Wang et al., 2023a; Jiang et al., 2023). For instance, in NLG,  Wang et al. (2023b) first introduces PandaLM, a fine-tuned LLM designed to evaluate pairwise texts based on several aspects. Similarly,  Li et al. (2023b) proposes Auto-J, an evaluator capable of handling a broader range of tasks and aspects, supporting both pointwise and pairwise evaluation settings. In image generation evaluation,  Xu et al. (2023a) develops ImageReward, an evaluator trained on a large-scale human annotation dataset. Using alignment and fidelity as reference dimensions, ImageReward provides an overall assessment. Despite showing good overall performance, coarse-grained evaluations lack the granularity needed to diagnose specific model deficiencies and often introduce aspect bias into the evaluation process.
2.2 Fine-grained Specific-task Evaluation
@@
To address the limitations of coarse-grained evaluation, recent studies have shifted toward fine-grained evaluation (Ye et al., 2023; Kim et al., 2024b; Ke et al., 2023; Kim et al., 2023; Li et al., 2023d; Ying et al., 2024; Bai et al., 2023), where MLLMs are fine-tuned on multi-aspect datasets to produce aspect-specific judgments. For example, in NLG,  Hu et al. (2024b) proposes Themis, an LLM trained on the GPT-4 annotated NLG-Eval corpus. In IU, LLaVA-Critic (Xiong et al., 2024) is the first fine-grained evaluator integrating diverse criteria, showing strong correlation with GPT-4o. For IG,  Xu et al. (2024) introduces VisionReward, a VQA-based evaluator that assesses image quality across fine-grained aspects. While these methods improve over coarse-grained evaluation by targeting specific aspects, they struggle with scalability across different tasks and aspects.  Liu et al. (2023b) further investigates aspect generalization through X-Eval, a two-stage learning framework that incorporates auxiliary aspects and demonstrates generalization in NLG. However, its reliance on predefined reference aspects and lack of open-source implementation restricts its reproducibility.
3 Methodology
@@
To develop UFEval, we need to construct a large-scale, multi-modal, and aspect-level evaluation dataset. However, existing datasets predominantly focus on overall quality assessment rather than fine-grained aspect evaluation, limiting the development of such evaluators. To address this gap, we construct FRABench through two main steps: (1) Evaluation Aspect Construction, and (2) Fine-grained Evaluation Dataset Construction. The following sections detail these two steps.
3.1 Evaluation Aspect Construction
3.1.1 Aspect Collection and Extension.
@@ 3.1 Evaluation Aspect Construction
3.1.2 Aspect Taxonomy Construction.
@@
Figure 1: An illustration of the evaluation pipeline. The pipeline consists of two steps: (1) Aspect Selection: First, appropriate aspects are selected from the TAs and UAs Trees based on task property and output modality. As illustrated, the question focuses on “cat image next to the orange dog picture”, while the image shows no cat adjacent to the orange dog. This discrepancy enables evaluation of the model’s ability to avoid hallucinations, specifically its performance on Context Inconsistency. Second, given the text output modality, aspects are selected from the Text Branch in the UAs Tree. (2) Evaluating: UFEval generates feedback and scores based on the input content and selected aspects.
  • Universal Aspects (UAs): Aspects in this subtree exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image) and serve as fundamental quality aspects across tasks.
  • Task-specific Aspects (TAs): Aspects in this subtree exhibit task dependence, as they primarily focus on task completion rather than output quality. They are typically closely tied to the task type. For example, engagingness for story generation or accuracy for mathematical reasoning.
For hierarchical construction within the two subtrees, we directly adopt existing aspect tree structures from previous studies (Hu et al., 2024a; Ghosh et al., 2023; Huang et al., 2023; DeAlcala et al., 2023; Zeng et al., 2023; Huang et al., 2025). For aspects lacking predefined hierarchies, we determine their positions within the tree based on their definitions—either assigning them to locations within existing structures or establishing them as leaf nodes of root node. Based on the aspect tree, the fine-grained evaluation process can be streamlined, as illustrated in Figure 1.
3.2 Fine-grained Evaluation Dataset Construction
Based on the aspect tree, we can construct fine-grained evaluation datasets for training and testing UFEval. Specifically, we construct FRABench—a large-scale, multi-modal, and aspect-level evaluation dataset. Following prior studies (Kim et al., 2024a; Ye et al., 2024) that demonstrate pointwise scoring is more susceptible to contextual bias and available for reward model training, we adopt a pairwise comparison evaluation method in FRABench.
3.2.1 FRABench Construction.
We first collect questions from the datasets corresponding to the 28 sub-tasks we gathered, and generate pairwise samples. Specifically, we first obtain 29.3K response pairs from public datasets and generate the remaining 30.1K pairs using different MLLMs (Details regarding the MLLMs used can be found in Appendix B.1). After obtaining the queries and response pairs, we assign aspects from our aspects tree. As a result, each pairwise sample is annotated with an average of 8 UAs and 3 TAs (Detailed information on aspect assignment is provided in Appendix B.4). We then generate evaluation labels via two approaches: (1) Human annotations: directly using three-aspect human-annotated scores from ImageRewardDB (Xu et al., 2023a), supplemented with feedback generated by GPT-4o; (2) GPT-4o annotations: due to the lack of human annotations for most aspects, we use GPT-4o to generate evaluation labels for all other pairwise samples. During the GPT-4o annotation process, we observed that GPT frequently incorporates response correctness into its assessment of UAs. To address this, we provide only the response, excluding the query, when evaluating UAs. All templates are provided in Appendix G. Moreover, to mitigate position bias (Ye et al., 2024), we balance the number of samples where response 1 is preferred over response 2 and vice versa by reversing the response positions in half of the surplus samples from the majority class and re-annotating them accordingly (Appendix B.2 presents detailed analysis about position bias of FRABench). Finally, we generate 325k fine-grained evaluation labels.
Tasks Dataset Sub-Tasks Split Size
Public
NLG $\blacksquare$ Auto-J Summarization, Creative Writing, Rewriting, etc. Train / FRA-ID 2.5k / 0.6k
@@ 3.2 Fine-grained Evaluation Dataset Construction
Generated
IU $\blacksquare$ TextVQA Text Reasoning Train / FRA-ID 2.1k / 0.4 k
@@
Figure 2: Pairwise data statistics of the FRAbench. “Public” indicates data derived from existing public datasets, “Generated” denotes data synthesized through our generation pipeline. Detailed counts of evaluation labels for each sub-task are in Appendix B.4.
3.2.2 Dataset Partitioning and Evaluator Training.
@@
Due to the lack of a large-scale aspect-level benchmark to verify UFEval’s aspect generalizable capability and support our arguments, we can only divide FRABench into training, In-Domain test set (FRA-ID), and Out-of-Domain test set (FRA-OOD). The training and FRA-ID consist of 18 randomly selected sub-tasks from four tasks, covering 22 UAs and 35 TAs, while FRA-OOD contains 10 unseen sub-tasks with 28 seen UAs and 27 unseen TAs. Finally, the training set contains 255.4K samples, FRA-ID contains 45.2K samples, and FRA-OOD contains 24.4K samples. The statistics of FRABench are presented in Figure 2, with more detailed analysis provided in Appendix B.
@@
An evaluator’s main job is to judge model responses like humans would, so their effectiveness depends on how closely their judgments match human opinions. To measure this, we create human-annotated evaluation datasets. Specifically, we extract partial samples from FRA-ID and FRA-OOD, and recruited three humans with master’s degrees for annotation. Finally, we retain 6.9K in-domain evaluation samples (FRA-ID-H) and 6.0K out-of-domain evaluation samples (FRA-OOD-H), respectively. Details of the human annotation process and annotation consistency are provided in Appendix E. When training UFEval, we use SFT to fine-tune Qwen2-VL-7B-Instruct on the training set. The detailed training configurations and methods are presented in Appendix C.1.
4 Experiments